3s + 1d = £3.50
2d + 1s =£3.25
What does S and D equal

Answer by Alan3354(69443)

3s + 1d = £3.50 --> 2d + 6s = 7
2d + 1s =£3.25
2d + 6s = 7
------------- subtract
0d - 5s = -3.75
s = 0.75
--------
2d + s = 3.25
2d = 2.5
d = 1.25